# Effective and Evasive Fuzz Testing-Driven Jailbreaking Attacks against LLMs

### v1

@@ Abstract
Large Language Models (LLMs) have excelled in various tasks but are still vulnerable to jailbreaking attacks, where attackers create jailbreak prompts to mislead the model to produce harmful or offensive content. Current jailbreak methods either rely heavily on manually crafted templates, which pose challenges in scalability and adaptability, or struggle to generate semantically coherent prompts, making them easy to detect. Additionally, most existing approaches involve lengthy prompts, leading to higher query costs. In this paper, to remedy these challenges, we introduce a novel jailbreaking attack framework, which is an automated, black-box jailbreaking attack framework that adapts the black-box fuzz testing approach with a series of customized designs. Instead of relying on manually crafted templates, our method starts with an empty seed pool, removing the need to search for any related jailbreaking templates. We also develop three novel question-dependent mutation strategies using an LLM helper to generate prompts that maintain semantic coherence while significantly reducing their length. Additionally, we implement a two-level judge module to accurately detect genuine successful jailbreaks.
We evaluated our method on 7 representative LLMs and compared it with 5 state-of-the-art jailbreaking attack strategies. For proprietary LLM APIs, such as GPT-3.5 turbo, GPT-4, and Gemini-Pro, our method achieves attack success rates of over 90%, 80%, and 74%, respectively, exceeding existing baselines by more than 60%. Additionally, our method can maintain high semantic coherence while significantly reducing the length of jailbreak prompts. When targeting GPT-4, our method can achieve over 78% attack success rate even with 100 tokens. Moreover, our method demonstrates transferability and is robust to state-of-the-art defenses. We will open-source our codes upon publication.
@@ 1 Introduction
Nowadays, large language models (LLMs) like ChatGPT [38] have experienced widespread deployment, showcasing exceptional capabilities in comprehending and generating human-like text. Currently, ChatGPT has over 100 million users, and its website receives 1.8 billion visits per month [^1] https://explodingtopics.com/blog/chatgpt-users. However, as large language models (LLMs) become more integrated into society, concerns about their security and safety have arisen. A significant worry is the potential misuse for malicious purposes. In a real-life incident in May 2023, a criminal was arrested for exploiting ChatGPT to create fake news [^2] https://www.cnbc.com/2023/05/09/chinese-police-arrest-man-who-allegedly-used-chatgpt-to-spread-fake-news.html. Moreover, a recent Microsoft study highlighted that a notable number of attackers are using LLMs to craft phishing emails and develop ransomware and malware [^3] https://www.msn.com/en-us/money/other/hackers-with-ai-are-harder-to-stop-microsoft-says/ar-AA1hK2eV. Apart from misuse, LLMs are susceptible to jailbreaking attacks [33, 48, 10, 6, 49, 29, 56, 53, 43]. Attackers aim to craft malicious prompts to mislead LLMs, bypass safety features, and generate responses with harmful, discriminatory, violent, or sensitive content. We present some jailbreaking attack samples in Figure 1.
@@
In this paper, we introduce a novel jailbreaking attack framework, which automatically generates concise, meaningful, and fluent jailbreak prompts. To enhance attack performance, we tackle the following challenges:
@@
C3. How to effectively judge the successful jailbreaking templates? To determine if a response is jailbroken, the judge model must evaluate both its harmfulness and its relevance to the given question. In our method, we utilize a two-level judge module to accurately identify true successful jailbreaks. The first module is a fine-tuned RoBERTa model that detects illegal content, and the second is a ChatGPT-based model that checks if the response matches the query and verifies the jailbreak status.
@@
  • We propose a fuzz-testing-driven jailbreaking attack framework, which can automatically generate jailbreak prompts with black-box access to the victim LLMs. Unlike existing works that rely on current jailbreaking template resources, our method starts the attack with an empty seed pool, removing the need to search for related jailbreaking templates and enhancing the practicality of the attack.
@@
  • Extensive experiments on LLaMA-2-7b-chat, Vicuna-7b-v1.3, Baichuan2-7b-chat, Guanaco-7B, GPT-3.5 Turbo, GPT-4, and Gemini-Pro LLM models verify the effectiveness and efficiency of our method compared to 5 state-of-the-art jailbreaking attacks. our method achieves a higher jailbreaking success rate than existing baselines by more than 60% across various LLMs, particularly when query tokens are shortened and defenses are in place. our method also demonstrates resistance to state-of-the-art jailbreaking defenses and showcases transferability.
@@ 2.2 Jailbreaking Attacks
Figure 2: Overview attack procedure of our method. our method features a fuzz testing-enabled jailbreaking attack. We initialize the attack with an empty seed pool, introduce three mutation strategies capable of generating templates with semantic coherence and limited prompt length, and design a two-level judge module to distinguish successful jailbreaks.
@@
Automatic Attacks. To enhance efficiency, automatic jailbreaking attacks have gained popularity. Deng et al. [11] introduced JailBreaker, a method that leverages an LLM to automatically generate effective jailbreak prompts. They curated and refined a specialized dataset of jailbreak prompts, trained a dedicated LLM for breaching chatbots, and incorporated a reward-ranked tuning step to boost the model’s ability to circumvent various LLM chatbot defenses. Similarly, Chao et al. [7] proposed PAIR, an approach that uses an attacker LLM to autonomously generate jailbreaks for a target LLM without human intervention. This method involves the attacker LLM iteratively querying the target LLM to refine and enhance the jailbreak. Additionally, Mehrotra et al. [35] introduced TAP, a method that employs an LLM to progressively refine potential attack prompts using tree-of-thoughts reasoning until a successful jailbreak is achieved. Before submitting prompts to the target, TAP evaluates and discards those unlikely to succeed, thereby effectively reducing the overall number of queries required.
In addition to relying on a helper model, attackers can also leverage genetic methods to automatically generate jailbreak prompts. Liu et al. [31] introduced AutoDAN, a novel jailbreak attack against aligned LLMs. AutoDAN utilizes a carefully designed hierarchical genetic algorithm to automatically generate stealthy jailbreak prompts. However, this method requires access to the LLM logits and should be considered as a gray-box attack. Lapid et al. [27] also utilized genetic procedures to generate suffixes for harmful questions. They start with a random token set as a seed suffix and employ token replacement as mutators. The selection of new seeds is rewarded based on how closely the target LLM’s response matches a desired pattern (e.g., “Sure, here is", followed by a prespecified answer to the harmful question). Recently, Yu et al. [55] proposed using fuzz testing to generate jailbreak prompts. They began with existing handcrafted templates as initial seeds and introduced five mutators to create new templates. Templates that lead the LLM to produce harmful content are selected as new seeds for further testing.
@@
  • Reliance on manual effort. While manual attacks can effectively discover stealthy jailbreak prompts, they often rely on individual LLM users crafting these prompts. This leads to challenges in scalability and adaptability.
  • Dependence on existing jailbreaking templates. Several existing jailbreak attacks rely on previously successful jailbreak prompts [11, 31]. These attacks use genetic algorithms or synonymous sentence conversion strategies to expand and enhance jailbreak prompts. However, once the original jailbreak prompts are mitigated, the efficacy of these attacks is consequently diminished. Moreover, finding a suitable template from numerous options can be costly and time-consuming, especially for new problems with limited references.
  • Oversight of prompt length. A prominent drawback in existing jailbreak attacks is the oversight of the length of the prompts. Typically, attackers generate longer jailbreak prompts to sustain the effectiveness of the jailbreak prompts [63, 7]. Given that the pricing structure for utilizing LLM APIs is influenced by factors like token or character length in both input and output, employing lengthy jailbreak prompts not only raises server alert concerns but also results in higher attack costs.
  • Oversight of semantic coherence. Many existing works overlook the semantic fluency and naturalness of generated prompts, resulting in outputs that are often uninterpretable to humans [63, 11]. Such non-natural prompts can be easily detected or nullified by existing jailbreaking defenses [42, 1].
  • Easily detected. Albeit effective for open-sourced LLMs or those deployed by small companies, many existing works lack robustness against more advanced and securely deployed LLMs, such as GPT-4 and Gemini-Pro. They also struggle against recently proposed jailbreaking defenses, such as [42, 17].
In this paper, we introduce a novel evasive jailbreaking attack, which autonomously generates jailbreak prompts without relying on any previously successful jailbreak prompts. Notably, our method is capable of producing short and natural jailbreak prompts but effectively circumventing the usage policy measures implemented by LLM Chatbots.
@@ 2.3 Jailbreaking Defenses
Output filtering-based defenses. These defenses evaluate whether the response is harmful and filter out malicious responses. Helbling et al. [16] leveraged the target LLM itself to assess whether a response is detrimental before its release. Li et al. [30] proposed generating harmless responses through an iterative process where each iteration feeds back into the LLM to produce and select response candidates based on a metric evaluating harmfulness and frequency. However, this defense can bring about a 4-fold increase in the inference time [60]. Inan et al. proposed Llama Guard [18], a supervised model that classifies prompt-response pairs as safe or unsafe. Xu et al. [52] fine-tuned a target LLM to reject specific jailbreak prompts and used this expert model to calibrate the outputs of the target LLM. Zeng et al. [58] proposed AutoDefense, which leveraged the response filtering ability of LLMs to identify unsafe responses triggered by jailbreak prompts. Kim et al. [24] proposed defending LLMs against jailbreaking attacks using “back translation". It starts by generating a response from a given input prompt using the target LLM and then uses another language model to infer an input prompt from this response, known as the backtranslated prompt. This new prompt can help uncover the true intent of the original prompt, and if the target LLM rejects the backtranslated prompt, the original prompt is also rejected.
@@
In this paper, we will evaluate whether our method is robust to existing jailbreaking defenses.
@@ 2.4 Fuzz Testing
We operates under the black-box fuzzing paradigm, whereby it does not access the source code or internal weights of the models under test. This process entails a series of strategic steps designed to uncover vulnerabilities: seed initialization, seed selection, mutation, and execution. Initially, seed initialization involves creating a starting input for the system, which could be either randomly generated or specially crafted to induce specific behaviors in the program. Subsequently, a seed is chosen from the pool, and this selection may be random or based on a heuristic designed to maximize coverage or explore new execution paths. The chosen seed then undergoes mutation to produce new, varied inputs. Finally, in the execution phase, these mutated inputs are run within the system. If an input causes the program to crash or uncovers a previously unexplored path, it is incorporated into the seed pool to facilitate future testing efforts.
In our method, we adapted the black-box fuzzing framework by designing various innovative components to address the challenges of generating jailbreaking templates.
@@ 3 Threat Model
The objective of the attack is to craft jailbreak prompts [^9] The jailbreak prompt is the final input to the LLM, created by replacing the placeholders (denotes as [INSERT PROMPT HERE]) in the jailbreaking template with the harmful question. for a collection of harmful or unethical questions that compel the target LLM to provide actual responses.

### v4

# Papillon: Efficient and Stealthy Fuzz Testing-Powered Jailbreaks for LLMs

@@ Abstract
Large Language Models (LLMs) have excelled in various tasks but are still vulnerable to jailbreaking attacks, where attackers create jailbreak prompts to mislead the model to produce harmful or offensive content. Current jailbreak methods either rely heavily on manually crafted templates, which pose challenges in scalability and adaptability, or struggle to generate semantically coherent prompts, making them easy to detect. Additionally, most existing approaches involve lengthy prompts, leading to higher query costs. In this paper, to remedy these challenges, we introduce a novel jailbreaking attack framework called Papillon, which is an automated, black-box jailbreaking attack framework that adapts the black-box fuzz testing approach with a series of customized designs. Instead of relying on manually crafted templates, Papillon starts with an empty seed pool, removing the need to search for any related jailbreaking templates. We also develop three novel question-dependent mutation strategies using an LLM helper to generate prompts that maintain semantic coherence while significantly reducing their length. Additionally, we implement a two-level judge module to accurately detect genuine successful jailbreaks.
We evaluated Papillon on 7 representative LLMs and compared it with 5 state-of-the-art jailbreaking attack strategies. For proprietary LLM APIs, such as GPT-3.5 turbo, GPT-4, and Gemini-Pro, Papillon achieves attack success rates of over 90%, 80%, and 74%, respectively, exceeding existing baselines by more than 60%. Additionally, Papillon can maintain high semantic coherence while significantly reducing the length of jailbreak prompts. When targeting GPT-4, Papillon can achieve over 78% attack success rate even with 100 tokens. Moreover, Papillon demonstrates transferability and is robust to state-of-the-art defenses. [^1] Our code is available at https://zenodo.org/records/14737139.
@@ 1 Introduction
Nowadays, large language models (LLMs) like ChatGPT [37] have experienced widespread deployment, showcasing exceptional capabilities in comprehending and generating human-like text. Currently, ChatGPT has over 100 million users, and its website receives 1.8 billion visits per month [^2] https://explodingtopics.com/blog/chatgpt-users. However, as large language models (LLMs) become more integrated into society, concerns about their security and safety have arisen. A significant worry is the potential misuse for malicious purposes. In a real-life incident in May 2023, a criminal was arrested for exploiting ChatGPT to create fake news [^3] https://www.cnbc.com/2023/05/09/chinese-police-arrest-man-who-allegedly-used-chatgpt-to-spread-fake-news.html. Moreover, a recent Microsoft study highlighted that a notable number of attackers are using LLMs to craft phishing emails and develop ransomware and malware [^4] https://www.msn.com/en-us/money/other/hackers-with-ai-are-harder-to-stop-microsoft-says/ar-AA1hK2eV. Apart from misuse, LLMs are susceptible to jailbreaking attacks [32, 47, 10, 6, 48, 28, 55, 52, 42, 8]. Attackers aim to craft malicious prompts to mislead LLMs, bypass safety features, and generate responses with harmful, discriminatory, violent, or sensitive content. We present some jailbreaking attack samples in Figure 1.
@@
In this paper, we introduce a novel jailbreaking attack framework, called Papillon, which automatically generates concise, meaningful, and fluent jailbreak prompts. To enhance attack performance, we tackle the following challenges:
@@
C3. How to effectively judge the successful jailbreaking templates? To determine if a response is jailbroken, the judge model must evaluate both its harmfulness and its relevance to the given question. In Papillon, we utilize a two-level judge module to accurately identify true successful jailbreaks. The first module is a fine-tuned RoBERTa model that detects illegal content, and the second is a ChatGPT-based model that checks if the response matches the query and verifies the jailbreak status.
@@
  • We propose a fuzz-testing-driven jailbreaking attack framework, Papillon, which can automatically generate jailbreak prompts with black-box access to the victim LLMs. Unlike existing works that rely on current jailbreaking template resources, Papillon starts the attack with an empty seed pool, removing the need to search for related jailbreaking templates and enhancing the practicality of the attack.
@@
  • Extensive experiments on LLaMA-2-7b-chat, Vicuna-7b-v1.3, Baichuan2-7b-chat, Guanaco-7B, GPT-3.5 Turbo, GPT-4, and Gemini-Pro LLM models verify the effectiveness and efficiency of Papillon compared to 5 state-of-the-art jailbreaking attacks. Papillon achieves a higher jailbreaking success rate than existing baselines by more than 60% across various LLMs, particularly when query tokens are shortened and defenses are in place. Papillon also demonstrates resistance to state-of-the-art jailbreaking defenses and showcases transferability.
@@ 2.2 Jailbreaking Attacks
Automatic Attacks. To enhance efficiency, automatic jailbreaking attacks have gained popularity. Deng et al. [11] introduced JailBreaker, a method that leverages an LLM to automatically generate effective jailbreak prompts. They curated and refined a specialized dataset of jailbreak prompts, trained a dedicated LLM for breaching chatbots, and incorporated a reward-ranked tuning step to boost the model’s ability to circumvent various LLM chatbot defenses. Similarly, Chao et al. [7] proposed PAIR, an approach that uses a static attacker LLM (fine-tuned using prior jailbreaking templates) to autonomously generate jailbreaks for a target LLM without human intervention. The attacker LLM iteratively queries the target LLM, using its responses to refine and enhance the prompts. However, its performance significantly declines against models with robust fine-tuning, such as Llama-2. Additionally, Mehrotra et al. [34] introduced TAP, a method that employs an LLM to progressively refine potential attack prompts using tree-of-thoughts reasoning until a successful jailbreak is achieved. Before submitting prompts to the target, TAP evaluates and discards those unlikely to succeed, thereby effectively reducing the overall number of queries required. Zeng et al. [57] drew inspiration from persuasion techniques in social science for LLM jailbreaking. This approach developed Persuasive Paraphrasers using examples under persuasion taxonomies to transform harmful queries into interpretable persuasive adversarial prompts (PAPs). However, it is vulnerable to advanced defenses and inadequate handling of multi-turn dialogues or complex persuasion scenarios.
In addition to relying on a helper model, attackers can also leverage genetic methods to automatically generate jailbreak prompts. Liu et al. [30] introduced AutoDAN, a novel jailbreak attack against aligned LLMs. AutoDAN utilizes a carefully designed hierarchical genetic algorithm to automatically generate stealthy jailbreak prompts. However, this method requires access to the LLM logits and should be considered as a gray-box attack. Lapid et al. [26] also utilized genetic procedures to generate suffixes for harmful questions. They start with a random token set as a seed suffix and employ token replacement as mutators. The selection of new seeds is rewarded based on how closely the target LLM’s response matches a desired pattern (e.g., “Sure, here is", followed by a prespecified answer to the harmful question). More recently, fuzzy techniques have been introduced to generate jailbreak prompts [52, 54]. These methods employ fuzz testing to automate the creation of jailbreak prompts. Starting with existing handcrafted templates as initial seeds, they apply various mutators to create new templates. Templates that lead the LLM to produce harmful content are selected as new seeds for further testing. While effective for open-source models, these techniques often fail when applied to advanced proprietary models like GPT-4 and Gemini-Pro.
@@
  • Reliance on manual effort. Manual attacks can discover stealthy jailbreak prompts, but are not scalable due to their reliance on individual LLM users. In addition, previous attacks require an initial seed pool of human-written input (e.g., role-playing-style jailbreaks) to generate jailbreak variations [52, 54]. The attack success rate is highly sensitive to the quality and diversity of the seed pool.
  • Dependence on static jailbreaking templates or models. Many existing attacks leverage previously successful jailbreak prompts [11, 30, 57], and use strategies such as genetic algorithms, synonymous sentence transformations, or adherence to persuasion guidelines to expand and enhance these prompts. If the originally selected jailbreak prompts are weak, the efficacy of these attacks will be consequently affected. Additionally, some methods leverage static attacker models to generate prompts [7], but finding a suitable template from numerous options can be costly and time-consuming, especially for new problems with limited references.
  • Prompt length. Existing attacks always generate lengthy jailbreak prompts to achieve successful attacks [62, 7]. As pricing models of most commercial LLM APIs are determined by token or character length, lengthy jailbreak prompts not only raise server alerts but also result in higher attack costs.
  • Semantic coherence. Many existing attacks overlook the semantic fluency and naturalness of the generated prompts, resulting in outputs unintelligible to humans [62, 11]. Such unnatural prompts can be easily detected or nullified by existing jailbreaking defenses [41, 1].
  • Vulnerability to defenses. Although being effective against open-sourced LLMs, many existing attacks can be easily thwarted by more advanced LLMs, such as GPT-4 and Gemini-Pro. They also struggle against recently proposed jailbreaking defenses, such as [41, 16].
In this paper, we introduce a novel evasive jailbreaking attack, named Papillon, which autonomously generates jailbreak prompts without relying on previously successful jailbreak prompts. Notably, Papillon is capable of producing short and natural jailbreak prompts that can effectively circumvent defense measures.
@@ 2.3 Jailbreaking Defenses
Output filtering-based defenses. These defenses evaluate whether the response is harmful and filter out malicious responses. Helbling et al. [15] leveraged the target LLM itself to assess whether a response is detrimental before its release. Li et al. [29] proposed generating harmless responses through an iterative process where each iteration feeds back into the LLM to produce and select response candidates based on a metric evaluating harmfulness and frequency. However, this defense can bring about a 4-fold increase in the inference time [60]. Inan et al. proposed Llama Guard [17], a supervised model that classifies prompt-response pairs as safe or unsafe. Xu et al. [51] fine-tuned a target LLM to reject specific jailbreak prompts and used this expert model to calibrate the outputs of the target LLM. Zeng et al. [58] proposed AutoDefense, which leveraged the response filtering ability of LLMs to identify unsafe responses triggered by jailbreak prompts. Kim et al. [23] proposed defending LLMs against jailbreaking attacks using “back translation". It starts by generating a response from a given input prompt using the target LLM and then uses another language model to infer an input prompt from this response, known as the back-translated prompt. This new prompt can help uncover the true intent of the original prompt, and if the target LLM rejects the back-translated prompt, the original prompt is also rejected.
Figure 2: Workflow of Papillon. Papillon features a fuzz testing-enabled jailbreaking attack with two main phases: template mutation/optimization and a two-level judge module. Papillon begins with an empty seed pool and operates in two phases: pre-jailbreak and final-jailbreak. In the pre-jailbreak phase, Role-play and Contextualization mutations are used to generate initial jailbreak attempts and populate the empty seed pool. The final-jailbreak phase builds on this by adding Expand mutation to the process. A two-level judge module evaluates LLM responses to identify successful jailbreaks.
@@
In this paper, we will evaluate whether Papillon is robust to existing jailbreaking defenses.
@@ 2.4 Fuzz Testing
Papillon operates under the black-box fuzzing paradigm, whereby it does not access the source code or internal weights of the models under test. This process entails a series of strategic steps designed to uncover vulnerabilities: seed initialization, seed selection, mutation, and execution. Initially, seed initialization involves creating a starting input for the system, which could be either randomly generated or specially crafted to induce specific behaviors in the program. Subsequently, a seed is chosen from the pool, and this selection may be random or based on a heuristic designed to maximize coverage or explore new execution paths. The chosen seed then undergoes mutation to produce new, varied inputs. Finally, in the execution phase, these mutated inputs are run within the system. If an input causes the program to crash or uncovers a previously unexplored path, it is incorporated into the seed pool to facilitate future testing efforts.
In Papillon, we adapted the black-box fuzzing framework by designing various innovative components to address the challenges of generating jailbreaking templates.
@@ 3 Threat Model
The objective of the attack is to craft jailbreak prompts [^10] The jailbreak prompt is the final input to the LLM, created by replacing the placeholders (denotes as [INSERT PROMPT HERE]) in the jailbreaking template with the harmful question. for a collection of harmful or unethical questions that compel the target LLM to provide actual responses. Specifically, our goals are: 1) To circumvent the safety mechanisms of the target LLM, ensuring that it responds to a harmful query instead of opting not to answer. 2) To ensure that the responses to these questions are not only correct and relevant but also contain the harmful content intended by the query.
4 Methodology of Papillon
In this section, we first present the general attack framework of Papillon and then describe its key components.
4.1 Framework of Papillon
Papillon is an automated black-box jailbreaking attack method based on fuzz testing, which encompasses a cyclical regimen of seed initialization, seed selection, mutation, and execution.
  • Seed initialization. The first step is to generate or select initial inputs (seeds) from the current pool. We confront a bifurcated decision: leveraging pre-existing, human-crafted jailbreaking templates [54] or initiating seeds stochastically [26]. Human-written templates, however, often lack scalability and adaptability, making it difficult and time-consuming to find a suitable one, especially for new questions with limited references. Conversely, randomly initialized jailbreaking templates can result in meaningless sequences, making it challenging to create semantically coherent jailbreaks. To overcome these limitations, Papillon initializes with an empty seed pool, eliminating the need for existing jailbreaking templates. To address the initial challenge of limited seed availability, we introduce a pre-jailbreak phase where each problem undergoes several initial jailbreak attempts (e.g., 10 times). Successful jailbreaks from this phase will bypass the subsequent attack procedure to reduce the query budget. Note that the pre-jailbreak queries were included in the total query count. Thus, the overall process in Papillon consists of two stages: the pre-jailbreak and final-jailbreak stages. This dual-phase strategy enhances the efficiency and effectiveness of Papillon.
  • Seed selection. Seed selection chooses specific seeds that are likely to achieve the attack goals to undergo mutation. There exist various seed selection strategies, such as random selection, Round Robin, UCB, and MCTS-Explore [54]. Follow [54], we use MCTS-Explore in Papillon by default.
  • Mutation. These selected seeds undergo systematic modification through mutation to create new jailbreaking templates. Traditional mutation strategies employed by fuzzers [^11] https://lcamtuf.coredump.cx/afl/ are primarily designed for binary or structured data. Directly applying these strategies to natural language inputs can lead to syntactically incorrect or semantically nonsensical inputs. In addition, various studies [30, 54] utilize specific mutation operators, such as “generate", “crossover", “expand", “shorten", and “rephrase", to generate mutations based on existing human-written jailbreaking templates. However, since these mutation methods are closely tied to the initial templates and are not specifically adapted to the harmful question, they often produce many redundant and ineffective mutations. In Papillon, we design three novel question-dependent mutation strategies capable of generating mutations with semantic meaning and limited token length by leveraging the power of an LLM helper. As a result, we can significantly improve attack performance and reduce attack costs.
  • Execution. Finally, the mutated seeds are evaluated using a judge model to distinguish between successful and unsuccessful jailbreaks. Successful jailbreaking templates are retained in the seed pool, while unsuccessful ones are discarded. To determine whether the response indicates a successful jailbreak, attackers can employ either human annotators [32, 56] or LLMs [31, 43, 46] for labeling assistance. However, using human annotators is not scalable and impractical for automatic fuzzing. Additionally, LLMs also suffer from inaccuracies. In Papillon, we implement a two-level automatic judge module to distinguish genuine successful jailbreaks, further decreasing query costs to victim LLMs.
Please refer to the Appendix for more detailed explanations of certain terminologies. The overview of Papillon is shown in Figure 2, and the overall algorithm is detailed in Algorithm 1.